OpenStudy (anonymous):

What is the volume of a sphere with a radius of 9? Its not V = 3053.63 so... idk A. 729pi units3 B. 648pi units3 C. 972pi units3

OpenStudy (anonymous):

The formula for the volume of a sphere is 4*pi*a^3/3

OpenStudy (anonymous):

If the radius is a, 4*pi*3^3/3

OpenStudy (anonymous):

In these problem, they didn't calculate pi .-.

OpenStudy (anonymous):

Pi is just 3.14 simply

OpenStudy (anonymous):

Interesting. Using the correct equation for volume, you get 3053.63, just as you did. If you decide to set pi = 1, you get 972.

OpenStudy (anonymous):

The equation to find the volume of a sphere is the following pi*radius squared (to the second power)*4 So, the equation is 3.14*4 to the second power (16)*4 Pretty sure that's the right equation o-o

OpenStudy (anonymous):

4(3.14)*(9^3)/3

OpenStudy (anonymous):

Oh, I see. They left is as "pi." Time for me to stop for the night.

OpenStudy (anonymous):

My answer for V is 3052.08

OpenStudy (anonymous):

Like I tried the normal way of solving it using the formula for a sphere and got 3053.63 Then I tried not calculating pi and got some number thats not a choice ._. and know idk D:

OpenStudy (anonymous):

9^3 = 729 so the answer is going to have 729 in it

OpenStudy (anonymous):

That should help

OpenStudy (anonymous):

Yeah, the answers are in the form "# Pi" So, don't put in a number for pi. Leave it as "pi."

OpenStudy (anonymous):

I was working backwards from the problem

OpenStudy (anonymous):

The asnwer was 972 guys o:but thanks for helping out tho! (:

OpenStudy (anonymous):

I am surprised by that... 972

OpenStudy (anonymous):

\[V = \frac{4}{3} \pi r^3 = \frac{4}{3} \pi *9^3= 972 \pi\] The answers are in this form. As written, "972pi" So, just combine all the other stuff. Leave pi as "pi"

OpenStudy (anonymous):

Oh I see what I did... didn't multiply by 4/3